OpenStudy (anonymous):

Which equation can be simplified to find the inverse of y=x^2-7?

OpenStudy (anonymous):

Possible answers....

OpenStudy (wwwhhhaaattt?):

What do you think it is?

OpenStudy (anonymous):

I'm not entirely sure. The unit that explained how to solve this equation was way back at the begging of the semester, so I'm kind of lost.

OpenStudy (wwwhhhaaattt?):

Ok, Try to solve it for me.

OpenStudy (anonymous):

You mean find the inverse?

OpenStudy (wwwhhhaaattt?):

Follow the problem.

OpenStudy (anonymous):

I'm sorry. I don't understand. :(

OpenStudy (wwwhhhaaattt?):

Yes, Find the Inverse

OpenStudy (anonymous):

Ok, one minute.

OpenStudy (anonymous):

y=x^2-7 Switch x & y: x=y^2-7 Solve for y: \[\sqrt{x}=\sqrt{y^2} -7\] \[\sqrt{x}+7=y\]

OpenStudy (wwwhhhaaattt?):

Ok, From that what can you use to get the answer?

OpenStudy (anonymous):

x=y^2-7 must be simplified to get the inverse of y=x^2-7 and that's the answer they want.

OpenStudy (anonymous):

Is that right @WWWHHHAAATTT? ?

OpenStudy (wwwhhhaaattt?):

Do you think it is?